OpenStudy (anonymous):

Solve P = 2a + 3s for s in terms of P and a.

OpenStudy (anonymous):

Given P = 2a + 3s subtract 2a from both sides P - 2a = 3s and divide both sides by 3 for our answer of 

OpenStudy (anonymous):

P = 2a + 3s 3s = P - 2a s = (P - 2a)/3